### v3

# The Relative Entropy of Expectation and Price

Abstract
As operators acting on the undetermined final settlement of a derivative security, expectation is linear but price is non-linear. When the market of underlying securities is incomplete, non-linearity emerges from the bid-offer around the mid price that accounts for the residual risks of the optimal funding and hedging strategy. At the extremes, non-linearity also arises from the embedded options on capital that are exercised upon default. In this essay, these convexities are quantified in an entropic risk metric that evaluates the strategic risks, which is realised as a cost with the introduction of bilateral margin. Price is then adjusted for market incompleteness and the risk of default caused by the exhaustion of capital.
In the complete market theory, price is derived from a martingale condition. In the incomplete market theory presented here, price is instead derived from a log-martingale condition:
$p=-\frac{1}{\alpha}\log\mathbb{E}\exp[-\alpha P]$ (1)
for the price $p$ and payoff $P$ of a funded and hedged derivative security, where the price measure $\mathbb{E}$ has minimum entropy relative to economic expectations, and the parameter $\alpha$ matches the risk aversion of the investor. This price principle is easily applied to standard models for market evolution, with applications considered here including model risk analysis, deep hedging and decentralised finance.
I Introduction
A derivative transaction exchanges the initial price $p$ for a commitment to return the contractual final price $p+dp$ at a later time. In determining its economic viability, the counterparties assess whether the trade enhances or degrades the overall performance of their investment portfolios, balancing expected return against net risk. Asymmetry between the expectations and investment objectives of participants then drives market evolution, following the dynamic that trades execute when parties agree on price but disagree on value.
Capital is deployed to finance the transaction, entailing economic consequences for the investor beyond the contractual terms of the derivative. In addition to the expected final settlements, valuation depends on the funding of initial and variation margin and the level of default protection they confer. The impact of margin is typically developed as a valuation adjustment to a price model derived from the hedging of contractual settlements. In this essay, an entropic margin model is instead placed at the centre of the pricing methodology, providing the risk counterbalance to expected return in the investment strategy. There are numerous benefits to this approach.
Investment strategy
Traditional strategies balance expected return against risk. Here, the risk metric is encoded in the entropic margin model as an additional settlement. Maximising the return on capital net of all settlements naturally embeds risk management.
Incentivising margin
In this approach, the investment strategy accounts for both the funding costs and risk management benefits of margin, and quantifies this in the derivative valuation. Margin thus becomes a key consideration for pricing.
Incomplete markets
Market equilibrium identifies a unique price model with minimum entropy relative to economic expectations, calibrated to available funding and hedging. In the process, this resolves the ambiguity of pricing in incomplete markets.
Market and model risk
Uniquely identifying price from expectation, market and model risks are distinguished respectively as sensitivities to observed prices and subjective economic assumptions. This enables a hierarchical approach that recognises market tiering.
Decentralised finance
The level of margin is controlled by a single parameter, and the link with funding settlements is transparent and programmatic. Implemented in the scripting of smart contracts for derivatives, this supports the transition to decentralisation.
Advocates of deep hedging replace risk neutralisation with risk optimisation as the guiding principle for pricing, as data-driven models see the complex correlations among economic variables as opportunity rather than hindrance. Convex risk optimisation is similarly used here to derive the price measure from the expectation measure, though the objectives differ. Model risk is quantified as the residual risk after funding and hedging, which requires for its definition a precise understanding of the relationship between expectation and price. Taking this idea to its logical conclusion, the convex risk metric is implemented as margin held against model risk. Optimisation of the strategic return net of all settlements then naturally embeds risk optimisation.
Figure 1: The contractual settlements of the trade.
Figure 2: The contractual and funding settlements of the trade.
Figure 3: The contractual, funding and default settlements of the trade.
Suppose that two counterparties deploy initial capital $c_{1}>0$ and $c_{2}>0$ to the transaction and deposit the proceeds in a liquid margin account with positive unit price $b>0$. The first counterparty buys the derivative from the second counterparty, and the capital deployed must be sufficient to enter the transaction:
$-c_{2}\leq p\leq c_{1}$ (2)
With the proceeds and unused capital deposited in the margin account, the final margin available to settle the derivative is $(c_{1}-p)(1+db/b)$ for the first counterparty and $(c_{2}+p)(1+db/b)$ for the second counterparty. This is sufficient to settle the contractual commitments of the derivative when:
$-(c_{1}-p)(1+db/b)\leq p+dp\leq(c_{2}+p)(1+db/b)$ (3)
Define the default trigger levels:
$x_{1}:={}$ $(c_{1}-p)(1+db/b)+(p+dp)$ (4)
[…]
$x_{2}:={}$ $(c_{2}+p)(1+db/b)-(p+dp)$
[…]
The first counterparty defaults when $x_{1}<0$ and the second counterparty defaults when $x_{2}<0$, and these two conditions are mutually exclusive. The actual final settlement of the derivative, net of default, is then:
$p+dp+(-x_{1})^{+}-(-x_{2})^{+}$ (5)
and the final capital of the counterparties is:
$c_{1}+dc_{1}$ $=(x_{1})^{+}-(-x_{2})^{+}$ (6)
$c_{2}+dc_{2}$ $=(x_{2})^{+}-(-x_{1})^{+}$
respectively. These expressions for the change in capital are rearranged to:
$d\frac{c_{1}}{b}$ $=\left(-\frac{c_{1}}{b}-d\frac{p}{b}\right)^{\!+}-\left(d\frac{p}{b}-\frac{c_{2}}{b}\right)^{\!+}+d\frac{p}{b}$ (7)
$d\frac{c_{2}}{b}$ $=\left(d\frac{p}{b}-\frac{c_{2}}{b}\right)^{\!+}-\left(-\frac{c_{1}}{b}-d\frac{p}{b}\right)^{\!+}-d\frac{p}{b}$
demonstrating that the discounted return on capital matches the discounted return on the derivative adjusted by a put option on $d(p/b)$ with strike $-c_{1}/b$, exercised when the first counterparty defaults, and a call option on $d(p/b)$ with strike $c_{2}/b$, exercised when the second counterparty defaults. When assessing the viability of the trade, it is these net returns that must be considered, accounting for funding costs and the possibility that the settling counterparty defaults on its commitments.
This simple framework for bilateral margin is used in the following to derive equilibrium pricing expressions that account for funding and default. It is an idealisation of real trading activity, and in particular does not account for the default protections provided by additional capital covering the net risks of the institution. Extensions to multilateral margin, which can be incorporated into the framework, are not considered here.
The objective is to create foundations for pricing that do not rely on unrealistic assumptions on market completeness or the effectiveness of the trading strategy, properly account for residual model risk, and establish consistency across diverse applications including algorithmic trading, derivative pricing and hedging, and the modelling of margin and capital. In developing this objective, the entropic risk metric has dual application: as the determinant of margin used to protect against default; and as the a priori measure of performance used to optimise the investment strategy.
II Entropic investment strategies
The objective of the investor is to optimise the return $dc$ on the investment strategy, a task that depends on an initial assessment of the final economy that is both subjective and indeterminate. This assessment is encapsulated in the expectation measure $\mathbb{E}$ of the investor. Define the corresponding probability, variance and covariance operators:
$\mathbb{P}[\mathcal{A}]$ $:=\mathbb{E}[1_{\mathcal{A}}]$ (8)
$\mathbb{V}[X]$ $:=\mathbb{E}[XX^{t}]-\mathbb{E}[X]\mathbb{E}[X]^{t}$
$\mathbb{V}[X,Y]$ $:=\mathbb{E}[XY^{t}]-\mathbb{E}[X]\mathbb{E}[Y]^{t}$
for the event $\mathcal{A}$ and the random vectors $X$ and $Y$. Investment performance is estimated using these operators, and the strategy is optimised accordingly.
II.1 Entropic risk metric
Maximising the mean $\mathbb{E}[dc]$ is an obvious strategic objective, though taken on its own this incentivises reckless investment and fails to reward effective risk management. Mean-variance optimisation extends the strategy by maximising the variance-adjusted mean:
$\mathbb{E}[dc]-\frac{1}{2}\alpha\mathbb{V}[dc]$ (9)
In this expression, the parameter $\alpha$ controls the risk aversion, and is interpreted as the Lagrange multiplier for the variance constraint in a strategy that maximises the mean for a fixed variance. Incorporating higher moments, entropic-risk optimisation further extends the strategy by maximising the entropy-adjusted mean:
$\mathbb{E}[\alpha][dc]:={}$ $\!-\frac{1}{\alpha}\log\mathbb{E}\exp[-\alpha\,dc]$ (10)
[…]
where $\mathbb{M}_{n}$ is the $n$th cumulant operator of the investor, so that $\mathbb{M}_{0}=0$, $\mathbb{M}_{1}=\mathbb{E}$ and $\mathbb{M}_{2}=\mathbb{V}$. The change of measure implicit in this definition reweights according to the sign and magnitude of the return, with properties that make it a useful metric of value-at-risk.
Figure 4: Entropy-adjusted mean as a function of risk aversion for three random variables. In each case, the variable is standardised by linear transformation to have zero mean and unit standard deviation. The entropy-adjusted mean is monotonic decreasing; it is unbounded for the normal variable, bounded below for the Poisson variable with ten expected jumps, and bounded above and below for the uniform variable.
Theorem II.1 (Entropic risk metric).
Let $dc$ be a random variable whose cumulant generating function in the probability measure $\mathbb{E}$ is defined. The entropy-adjusted mean:
$\mathbb{E}[\alpha][dc]:=-\frac{1}{\alpha}\log\mathbb{E}\exp[-\alpha\,dc]$ (11)
is monotonic decreasing as a function of the parameter $\alpha>0$ and has limits:
$\mathbb{E}[0][dc]$ $=\mathbb{E}[dc]$ (12)
$\mathbb{E}[\infty][dc]$ $=\operatorname*{ess\,inf}[dc]$
Let $\bar{\mathbb{E}}$ be an equivalent probability measure with strictly positive Radon-Nikodym weight $d\bar{\mathbb{E}}/d\mathbb{E}$. Define the entropy of $\bar{\mathbb{E}}$ relative to $\mathbb{E}$ by:
$S[\bar{\mathbb{E}}|\mathbb{E}]:=\mathbb{E}\!\left[\frac{d\bar{\mathbb{E}}}{d\mathbb{E}}\log\!\left[\frac{d\bar{\mathbb{E}}}{d\mathbb{E}}\right]\right]$ (13)
Relative entropy is a metric of distance between equivalent probability measures: it is positive, $S[\bar{\mathbb{E}}|\mathbb{E}]\geq 0$, and is zero when $\bar{\mathbb{E}}=\mathbb{E}$. Let $\mathbb{E}^{\alpha}$ be the equivalent probability measure with Radon-Nikodym weight $d\mathbb{E}^{\alpha}/d\mathbb{E}$ given by:
$\frac{d\mathbb{E}^{\alpha}}{d\mathbb{E}\hfill}:=\frac{\exp[-\alpha\,dc]}{\mathbb{E}[\exp[-\alpha\,dc]]}$ (14)
The entropy-adjusted mean $\mathbb{E}[\alpha][dc]$ equals the mean $\bar{\mathbb{E}}[dc]$ adjusted by relative entropy:
$\mathbb{E}[\alpha][dc]=\bar{\mathbb{E}}[dc]+\frac{1}{\alpha}(S[\bar{\mathbb{E}}|\mathbb{E}]-S[\bar{\mathbb{E}}|\mathbb{E}^{\alpha}])$ (15)
Instances of this equality for $\bar{\mathbb{E}}=\mathbb{E}$ and $\bar{\mathbb{E}}=\mathbb{E}^{\alpha}$ generate bounds for the entropy-adjusted mean:
[…]
Proof.
These are standard results from information theory, and only the key elements of their proofs are presented here. Monotonicity is a consequence of Hölder’s inequality:
$\mathbb{E}[\exp[-\alpha_{1}dc]]^{1/\alpha_{1}}\leq\mathbb{E}[\exp[-\alpha_{2}dc]]^{1/\alpha_{2}}$ (17)
for $0<\alpha_{1}<\alpha_{2}$, so that $\mathbb{E}[\alpha_{1}][dc]\geq\mathbb{E}[\alpha_{2}][dc]$. The first limit follows from the expansion of the cumulant generating function. For the second limit, Markov’s inequality:
$\mathbb{P}[dc\leq\gamma]\leq\frac{\mathbb{E}[\exp[-\alpha\,dc]]}{\exp[-\alpha\gamma]}$ (18)
for the parameters $\alpha>0$ and $\gamma>\operatorname*{ess\,inf}[dc]$ (so that $\mathbb{P}[dc\leq\gamma]>0$) is rearranged to:
$\mathbb{E}[\alpha][dc]\leq\gamma-\frac{1}{\alpha}\log[\mathbb{P}[dc\leq\gamma]]$ (19)
From this, taking the limit $\alpha\to\infty$ followed by taking the limit $\gamma\to\operatorname*{ess\,inf}[dc]$ derives $\mathbb{E}[\infty][dc]\leq\operatorname*{ess\,inf}[dc]$. The other side of the target inequality follows from the almost certainty of $\exp[-\alpha\,dc]\leq\exp[-\alpha\operatorname*{ess\,inf}[dc]]$, so that $\mathbb{E}[\alpha][dc]\geq\operatorname*{ess\,inf}[dc]$.
Gibb’s theorem states that $S[\bar{\mathbb{E}}|\mathbb{E}]\geq 0$, which follows from the observation that $x\log[x]\geq x-1$ for $x>0$. The remaining result is a presentation of Donsker and Varadhan’s variational formula. Let $\bar{m}:=d\bar{\mathbb{E}}/d\mathbb{E}$ and $m_{\alpha}:=d\mathbb{E}^{\alpha}/d\mathbb{E}$ be the Radon-Nikodym weights. Then:
$S[\bar{\mathbb{E}}$ $|\mathbb{E}]-S[\bar{\mathbb{E}}|\mathbb{E}^{\alpha}]$ (20)
$=\mathbb{E}[\bar{m}\log[\bar{m}]]-\mathbb{E}^{\alpha}[(\bar{m}/m_{\alpha})\log[\bar{m}/m_{\alpha}]]$
$=\bar{\mathbb{E}}[\log[\bar{m}]]-(\bar{\mathbb{E}}[\log[\bar{m}]]-\bar{\mathbb{E}}[\log[m_{\alpha}]])$
$=\bar{\mathbb{E}}[-\alpha\,dc-\log\mathbb{E}\exp[-\alpha\,dc]]$
$=-\alpha(\bar{\mathbb{E}}[dc]-\mathbb{E}[\alpha][dc])$
relating the entropy-adjusted mean to the relative entropies of the measures. ∎
The lever $\alpha>0$ thus dials the entropy-adjusted mean $\mathbb{E}[\alpha][dc]$ across the span of possible values for $dc$ between $\operatorname*{ess\,inf}[dc]$ and $\mathbb{E}[dc]$ and no further. Entropic risk is, in this sense, more targeted than variance adjustment as a moments-based metric of value-at-risk. The equivalent measure $\mathbb{E}^{\alpha}$ plays a central role in this result, and the gap between the mean $\mathbb{E}[dc]$ and the entropy-adjusted mean $\mathbb{E}[\alpha][dc]$ is proportional to the distance between the measure $\mathbb{E}$ and the measure $\mathbb{E}^{\alpha}$, as defined by their relative entropy.
II.2 Optimising entropic risk
The investor has a choice of market strategies whose returns over the investment period are the components of the return vector $dq$. The return on the investment portfolio whose weights are the components of the weight vector $\omega$ is then $dc:=\omega\cdot dq$, the weighted sum of the returns on the market strategies.
Mean-variance and entropic-risk optimisation are the basis for investment strategies, leading in each case to equilibrium models of pricing and hedging. There are many similarities between these strategies, but also some critical differences. Mean-variance optimisation does not account for non-Gaussian features in the return, and the strategy may need to be moderated to avoid arbitrage. In contrast, entropic-risk optimisation incorporates all the cumulants of the return, and naturally leads to arbitrage-free equilibrium pricing.
Introduce the families of equivalent measures $\mathbb{E}^{\phi}$ and $\mathbb{E}^{\phi,\zeta}$ with Radon-Nikodym weights:
$\frac{d\mathbb{E}^{\phi}}{d\mathbb{E}\hfill}$ $:=\frac{\exp[-\phi\cdot dq]}{\mathbb{E}[\exp[-\phi\cdot dq]}$ (21)
$\frac{d\mathbb{E}^{\phi,\zeta}}{d\mathbb{E}\hfill}$ $:=\frac{\exp[-\phi\cdot dq-\frac{1}{2}dq\cdot\zeta\,dq]}{\mathbb{E}[\exp[-\phi\cdot dq-\frac{1}{2}dq\cdot\zeta\,dq]]}$
parametrised by the vector $\phi$ and the positive-definite matrix $\zeta$. Price measures that are equivalent to the expectation measure are identified within this collection from the equilibrium of the entropic-risk optimisation strategy. The parameter $\phi$ arises from the change of measure in the definition of the entropic risk metric, which depends on the existence of the cumulant generating function. If this does not exist, the quadratic-regularised measure with parameter $\zeta$ is used instead. Regularisation is removed in the limit $\zeta\to 0$; if the strategy does not survive this limit, then regularisation, which moderates the impact of large events, is retained as a parameter of the model. In the following, the cumulant generating function is assumed to exist, implying the use of quadratic regularisation wherever necessary.
In the entropic-risk optimisation strategy, the investor seeks the optimal portfolio $\omega$ of market strategies that maximises the entropy-adjusted mean:
$\mathbb{E}[\alpha][\omega\cdot dq]:=-\frac{1}{\alpha}\log\mathbb{E}\exp[-\alpha\omega\cdot dq]$ (22)
Risk aversion is controlled by the parameter $\alpha$ in the range $0\leq\alpha\leq\infty$, and entropic risk is monotonic in this range:
[…]
Large values of $\alpha$ thus imply that the investor takes more account of negative outcomes in the strategic objective, corresponding to a lower risk appetite. The optimal strategy is the stationary point of entropic risk $\omega=(1/\alpha)\phi$ where the unit optimal portfolio $\phi$ satisfies the calibration condition:
$0=\mathbb{E}^{\phi}[dq]$ (24)
As a practical observation, the Newton-Raphson scheme:
$\phi^{(0)}$ $=0$ (25)
$\phi^{(n+1)}$ $=\phi^{(n)}+\mathbb{V}^{\phi^{(n)}}[dq]^{-1}\mathbb{E}^{\phi^{(n)}}[dq]$
calibrates the unit optimal portfolio to the calibration condition when the variance is invertible. This portfolio is a property of the market return, generating the price measure $\mathbb{E}^{\phi}$ equivalent to the expectation measure $\mathbb{E}$.
The optimal strategy is constructed on the assumption that the investor has expertise in all the available markets. In practice, markets are tiered and investors specialise in specific market sectors, and the optimal strategy across all sectors emerges as the net strategy of multiple specialist investors. This arrangement is conveniently captured in the entropic risk metric.
Suppose that the market return vector decomposes as a direct sum $dq=dq_{1}\oplus\cdots\oplus dq_{n}$ where $dq_{i}$ is the return vector for the $i$th market tier. The optimal portfolio is similarly decomposed as a direct sum $\phi=\phi_{1}\oplus\cdots\oplus\phi_{n}$ where $\phi_{i}$ is the portfolio in the $i$th market tier, and is solved from the calibration conditions:
$0=\mathbb{E}^{\phi_{1}\oplus\cdots\oplus\phi_{n}}[dq_{i}]$ (26)
for $i=1,\ldots,n$. This provides $n$ conditions for the $n$ components of the optimal portfolio. While these calibration conditions need to be solved simultaneously, market tiering is exploited to incrementally build the optimal portfolio from the independent contributions of sector investors, localising at each tier the risk management activity for the corresponding market expert. In this perspective, there are $n(n+1)/2$ calibration conditions for $n$ sub-markets:
00 0 $=\mathbb{E}^{\phi_{1}^{1}}[dq_{1}]$ (27)
00 0 $=\mathbb{E}^{\phi_{1}^{2}\oplus\phi_{2}^{2}}[dq_{1}]=\mathbb{E}^{\phi_{1}^{2}\oplus\phi_{2}^{2}}[dq_{2}]$
$\vdots$
00 0 $=\mathbb{E}^{\phi_{1}^{n}\oplus\cdots\oplus\phi_{n}^{n}}[dq_{1}]=\cdots=\mathbb{E}^{\phi_{1}^{n}\oplus\cdots\oplus\phi_{n}^{n}}[dq_{n}]$
where the calibration conditions in the $i$th row maximise the entropy-adjusted mean of the $i$th optimal portfolio $\phi^{i}=\phi_{1}^{i}\oplus\cdots\oplus\phi_{i}^{i}$ constituted in the lowest $i$ market tiers only. The $i$th investor then takes responsibility for determining the primary component $\psi_{i}$ of the $i$th optimal portfolio and its delta $\delta^{i}=\delta_{1}^{i}\oplus\cdots\oplus\delta_{i-1}^{i}$ to the $(i-1)$th optimal portfolio in the underlying sub-market. Stacking these deltas for $i=1,\ldots,n$ generates the optimal portfolio of the full market.
Theorem II.2 (Optimising entropic risk).
Consider a market of investment strategies whose return vector decomposes as a direct sum $dq=dq_{1}\oplus\cdots\oplus dq_{n}$ where $dq_{i}$ is the return vector for the $i$th tier of the market. For each $i=1,\ldots,n$, define the $i$th sub-market with return vector $dq^{i}=dq_{1}\oplus\cdots\oplus dq_{i}$ and let $\phi^{i}=\phi_{1}^{i}\oplus\cdots\oplus\phi_{i}^{i}$ be a portfolio in this sub-market. Use the invertible linear change of variables:
$\begin{array}[]{lr}\begin{array}[]{rl}\psi_{i}&=\phi_{i}^{i}\\ \delta_{j}^{i}&=\phi_{j}^{i-1}-\phi_{j}^{i}\end{array}&\qquad\phi_{j}^{i}=\psi_{j}-(\delta_{j}^{j+1}+\cdots+\delta_{j}^{i})\end{array}$ (28)
to represent the delta from the portfolio $\phi^{i-1}$ to the portfolio $\phi^{i}$ as the primary portfolio $\psi_{i}$ in the leading market tier minus the hedge portfolio $\delta_{1}^{i}\oplus\cdots\oplus\delta_{i-1}^{i}$ in the underlying market tiers.
Let $\mathbb{E}$ be a probability measure. For each $i=0,\ldots,n$, define the equivalent probability measure $\mathbb{E}^{i}$ with Radon-Nikodym weight:
$\frac{d\mathbb{E}^{i}}{d\mathbb{E}\hfill}:=\frac{\exp[-(\phi_{1}^{i}\cdot dq_{1}+\cdots+\phi_{i}^{i}\cdot dq_{i})]}{\mathbb{E}[\exp[-(\phi_{1}^{i}\cdot dq_{1}+\cdots+\phi_{i}^{i}\cdot dq_{i})]]}$ (29)
and define the $i$th entropy-adjusted mean:
$\mu^{i}:=-\log\mathbb{E}\exp[-(\phi_{1}^{i}\cdot dq_{1}+\cdots+\phi_{i}^{i}\cdot dq_{i})]$ (30)
At the $i$th tier, the portfolio $\phi^{i}$ maximises the entropy-adjusted mean $\mu^{i}$ when the primary portfolio $\psi_{i}$ and the hedge portfolio $\delta_{1}^{i}\oplus\cdots\oplus\delta_{i-1}^{i}$ are calibrated to the hedge conditions:
$0=$ (31)
$\mathbb{E}^{i-1}[dq_{j}\exp[-(\psi_{i}\cdot dq_{i}-(\delta_{1}^{i}\cdot dq_{1}+\cdots+\delta_{i-1}^{i}\cdot dq_{i-1}))]]$
for $j=1,\ldots,i$. Starting with the probability measure $\mathbb{E}^{0}=\mathbb{E}$, these calibrations are iterated to obtain the unit optimal portfolio $\phi^{n}$ in the full market, with its associated equivalent probability measure $\mathbb{E}^{n}$ calibrated to the calibration conditions for the market strategies.
Proof.
This result follows from the measure-change relationship:
$\log{}$ $\mathbb{E}^{i-1}\exp[-(\psi_{i}\cdot dq_{i}-(\delta_{1}^{i}\cdot dq_{1}+\cdots+\delta_{i-1}^{i}\cdot dq_{i-1}))]$
[…]
$-\log\mathbb{E}\exp[-(\phi_{1}^{i-1}\cdot dq_{1}+\cdots+\phi_{i-1}^{i-1}\cdot dq_{i-1})]$
between the entropy-adjusted means. ∎
Since expanding the market can only increase the opportunities for the investor, the maximum entropy-adjusted mean increases, $\mu^{i}\geq\mu^{i-1}$, as new tiers are introduced. Furthermore, the $i$th tier introduces no additional advantage for the strategy if the optimal portfolios $\phi^{i}$ and $\phi^{i-1}$ satisfy the indifference condition $\mu^{i}=\mu^{i-1}$. This is adopted as the equilibrium condition for pricing derivative strategies that have no intrinsic value beyond their contractual link with underlying strategies.
III Entropic pricing and hedging
The market of underlying securities accessible to the investor performs the two key functions of funding and hedging for the derivative security. Fix a finite investment horizon, and assume the market is tiered into a set of underlying securities with initial price vector $q$ and final price vector $q+dq$ and a derivative security with initial price $p$ and final price $p+dp$. The investor quantifies a priori the performance of the investment strategy using an initial assessment of the final economic state, encapsulated in the expectation measure $\mathbb{E}$ that maps variables at final time $t+dt$ to their expectations at initial time $t$. Following the strategy developed in the previous section, the optimal portfolio then maximises the entropy-adjusted mean return at a specified risk aversion $\alpha$ for the investor.
As summarised in Figure 5, there are two self-funded investment strategies to consider: one that includes only underlying securities; and one that also includes the derivative security. In the first strategy, the portfolio $\omega$ of underlying securities is optimal when it maximises the entropy-adjusted mean return $\mathbb{E}[\alpha][\omega\cdot dq]$. The optimal portfolio thus satisfies the calibration condition:
$0=\mathbb{E}[(dq-qr\,dt)\exp[-\alpha\omega\cdot dq]]$ (33)
where $r\,dt$ is the Lagrange multiplier used to implement the self-funding condition $\omega\cdot q=0$. Adjusting for the opportunity and risk that adding the derivative security presents, in the second strategy the optimal portfolio $\omega$ is modified by the hedge portfolio $\delta$, and optimality is restored in the combined portfolio when it maximises the entropy-adjusted mean return $\mathbb{E}[\alpha][dp+(\omega-\delta)\cdot dq]$. The hedge portfolio thus satisfies the hedge condition:
$0=\mathbb{E}[(dq-q(r+\alpha s)\,dt)\exp[-\alpha(dp+(\omega-\delta)\cdot dq)]]$ (34)
where $(r+\alpha s)\,dt$ is the Lagrange multiplier used to implement the self-funding condition $\delta\cdot q=p$.
Returns from correlated changes in the derivative and underlying prices are cancelled by hedging, with the residual return on the combined portfolio maximising the entropy-adjusted mean. When its price additionally satisfies the price condition:
$\mathbb{E}[\alpha][\omega\cdot dq]=\mathbb{E}[\alpha][dp+(\omega-\delta)\cdot dq]$ (35)
the derivative security neither advantages nor disadvantages the investment strategy.
Entropic Pricing and Hedging
Fix a finite investment horizon and let $\mathbb{E}$ be the expectation measure mapping variables at the final time $t+dt$ to their expectations at the initial time $t$. Assume the market is tiered into a set of underlying securities with initial price vector $q$ and final price vector $q+dq$ and a derivative security with initial price $p$ and final price $p+dp$. Consider two self-funded strategies: one comprising underlying securities only; and one that also includes the derivative security.
Underlying only: In this strategy, the investment comprises an underlying portfolio $\omega$ satisfying the initial condition: $0=\omega\cdot q$ (36) so that the investment is self-funded. The return on the investment has entropy-adjusted mean: $\mathbb{E}[\alpha]$ $[\omega\cdot dq]=$ (37) $-\frac{1}{\alpha}\log\mathbb{E}\exp[-\alpha\omega\cdot dq]$ at the risk aversion $\alpha>0$. The optimal portfolio $\omega$ that maximises this satisfies the calibration condition: $0=\mathbb{E}[(dq-qr\,dt)\exp[-\alpha\omega\cdot dq]]$ (38) where $r\,dt$ is the Lagrange multiplier used to implement the self-funding condition. Underlying and derivative: In this strategy, the investment comprises the derivative security and an underlying portfolio $(\omega-\delta)$ satisfying the initial condition: $0=p+(\omega-\delta)\cdot q$ (39) so that the investment is self-funded. The return on the investment has entropy-adjusted mean: $\mathbb{E}[\alpha]$ $[dp+(\omega-\delta)\cdot dq]=$ (40) $-\frac{1}{\alpha}\log\mathbb{E}\exp[-\alpha(dp+(\omega-\delta)\cdot dq)]$ at the risk aversion $\alpha>0$. The hedge portfolio $\delta$ that maximises this satisfies the hedge condition: $0=\mathbb{E}[(dq-q(r+\alpha s)\,dt)\exp[-\alpha(dp+(\omega-\delta)\cdot dq)]]$ (41) where $(r+\alpha s)\,dt$ is the Lagrange multiplier used to implement the self-funding condition.
These two strategies offer different entropy-adjusted mean returns to the investor. When the derivative security has no intrinsic value beyond its contractual link with the underlying securities, the efficient market equilibrates to the indifferent state wherein these entropy-adjusted means align, leading to the price condition: $\mathbb{E}[\alpha][\omega\cdot dq]=\mathbb{E}[\alpha][dp+(\omega-\delta)\cdot dq]$ (42) The calibration condition is solved to derive the optimal portfolio $\omega$ of underlying securities. The hedge and price conditions are then solved simultaneously to derive the initial price $p$ from the final price $p+dp$ of the derivative security and the hedge portfolio $\delta$ of underlying securities that re-establishes optimality for the combined portfolio.
Figure 5: In entropic pricing, economic principles are established that determine the optimal portfolio of underlying securities and the hedge portfolio and price of the derivative security, based on the maximisation of the entropy-adjusted mean return.
Refer to caption Refer to caption
$\rho=-0.5$ $\rho=-0.5$
Refer to caption Refer to caption
$\rho=0$ $\rho=0$
Refer to caption Refer to caption
$\rho=0.5$ $\rho=0.5$
Figure 6: Discounting is an emergent feature in entropic pricing, where the funding rate is implied as the Lagrange multiplier that implements the self-funding condition. In this example, two underlying securities are modelled with lognormal final prices $q+dq=q(1+\mu\,dt)\exp[\sigma\sqrt{dt}\,z-\sigma^{2}dt/2]$ in the expectation measure, using standard normal variables $z$ that have correlation $\rho$. The model for the second underlying security is fixed with $q_{2}=1$, $\mu_{2}=5\%$ and $\sigma_{2}=20\%$. The model for the first underlying security uses $q_{1}=1$, and the graphs in each row plot the implied funding rate $r$ and unit optimal portfolio $\phi_{1}$ as $\mu_{1}$ and $\sigma_{1}$ are varied. The correlation is $\rho=-0.5$ in the first row, $\rho=0$ in the second row, and $\rho=0.5$ in the third row, and the investment horizon is set to $dt=1$.
      When the volatility $\sigma_{1}$ of the first underlying security is small, the implied funding rate $r$ aligns with the mean rate $\mu_{1}$, as in this case the underlying security approximates a funding security with predictable return. When the volatilities $\sigma_{1}$ and $\sigma_{2}$ of the underlying securities are close to each other, the implied funding rate $r$ is approximately the average of the two mean rates $\mu_{1}$ and $\mu_{2}$. The mean rates of the underlying securities thus contribute to the implied funding rate, which is essentially the predictability-weighted average of the mean rates.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 7: Three examples of entropic hedging and pricing for a call option on an underlying security with return $dq$ modelled as a normal variable, a Poisson variable with ten expected jumps, and a uniform variable in the price measure. In each case, the return is standardised by linear transformation to have zero mean and unit standard deviation, and the derivative payoff for strike $k$ is $P:=(dq-k)^{+}$. Risk aversion is fixed at one, $\alpha=1$, and the implied funding rate is set to zero, $r=0$. The graphs plot the delta and price per unit notional as a function of the strike for a range of notionals $\lambda$, highlighting the central plot for the mid-pricing case $\lambda=0$. Residual risk after hedging creates a bid-offer spread around the mid price and skews the optimal hedge ratio for positions with large notional.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 8: Three examples of entropic hedging and pricing for a digital option on an underlying security with return $dq$ modelled as a normal variable, a Poisson variable with ten expected jumps, and a uniform variable in the price measure. In each case, the return is standardised by linear transformation to have zero mean and unit standard deviation, and the derivative payoff for strike $k$ is $P:=1_{dq>k}$. Risk aversion is fixed at one, $\alpha=1$, and the implied funding rate is set to zero, $r=0$. The graphs plot the delta and price per unit notional as a function of the strike for a range of notionals $\lambda$, highlighting the central plot for the mid-pricing case $\lambda=0$. Residual risk after hedging creates a bid-offer spread around the mid price and skews the optimal hedge ratio for positions with large notional.
Calibration generates the implied funding rate $r$ and the price measure $\mathbb{E}^{\phi}$ equivalent to the expectation measure $\mathbb{E}$ with Radon-Nikodym weight:
$\frac{d\mathbb{E}^{\phi}}{d\mathbb{E}\hfill}:=\frac{\displaystyle\exp[-\phi\cdot dq]}{\displaystyle\mathbb{E}[\exp[-\phi\cdot dq]]}$ (43)
where the unit optimal portfolio $\phi$ satisfies the self-funding condition $\phi\cdot q=0$ and the calibration condition:
$qr\,dt=\mathbb{E}^{\phi}[dq]$ (44)
All investments are self-funded within the underlying market, and this model does not assume the existence of a risk-free funding rate for investments that are not self-funded. The implied funding rate defined here originates in the optimisation as the Lagrange multiplier for the self-funding condition.
Let $d\bar{q}:=dq-qr\,dt$ be the underlying return centralised in the price measure and let $P:=p+dp$ be the final derivative price. Using the price measure, the hedge and price conditions are re-expressed as:
00 0 $=\mathbb{E}^{\phi}[(d\bar{q}-\alpha qs\,dt)\exp[-\alpha(P-\delta\cdot d\bar{q})]]$ (45)
$p$ $=\frac{1}{1+r\,dt}\mathbb{E}^{\phi}[\alpha][P-\delta\cdot d\bar{q}]$
These conditions express a non-linear relationship between the initial and final derivative prices. By definition, the derivative security with notional $\lambda$ has final price $\lambda P$. The corresponding initial price $p_{\lambda}$ and hedge portfolio $\delta_{\lambda}$ per unit notional and the implied funding spread $s_{\lambda}$ are determined from the hedge and price conditions:
00 0 $=\mathbb{E}^{\phi}[(d\bar{q}-\alpha qs_{\lambda}\,dt)\exp[-\lambda\alpha(P-\delta_{\lambda}\cdot d\bar{q})]]$ (46)
$p_{\lambda}$ $=\frac{1}{1+r\,dt}\mathbb{E}^{\phi}[\lambda\alpha][P-\delta_{\lambda}\cdot d\bar{q}]$
and the self-funding condition. Scaling the derivative notional is thus equivalent to scaling the risk aversion.
When $\alpha$ is small, the hedge and price conditions are expanded to discover the solutions:
$s$ $=\frac{\displaystyle\mathbb{E}^{\phi}[P]/(1+r\,dt)-\mathbb{V}^{\phi}[dq]^{-1}\mathbb{V}^{\phi}[dq,P]\cdot q}{\displaystyle\mathbb{V}^{\phi}[dq]^{-1}q\,dt\cdot q}+O[\alpha]$
$\delta$ $=\mathbb{V}^{\phi}[dq]^{-1}(\mathbb{V}^{\phi}[dq,P]+qs\,dt)+O[\alpha]$ (47)
$p$ $=\frac{1}{1+r\,dt}\left(\mathbb{E}^{\phi}[P]-\frac{1}{2}\alpha\mathbb{V}^{\phi}[P-\delta\cdot dq]\right)+O[\alpha^{2}]$
This expansion in terms of $\alpha$ approximates the price when the risk appetite of the investor is large or when the derivative security has small notional. In the marginal market, the model thus implies a limit order book for the derivative security with mid price and bid-offer spread:
$p_{\text{mid}}$ $=\mathbb{E}^{\phi}[P]/(1+r\,dt)$ (48)
$p_{\text{bid-offer}}$ $=\alpha\mathbb{V}^{\phi}[P-\delta\cdot dq]/(1+r\,dt)$
The discounted mid price of the derivative security is a martingale in the price measure and the bid-offer spread prices incompleteness of the underlying market. Higher-order adjustments then accommodate the extended features of the price distributions.
This model is extensive in scope, consistently encompassing arbitrary economic assumptions, predictable and unpredictable funding, complete and incomplete hedge markets, continuous and discontinuous evolution, continuous and discrete hedging, and accounting for the risk appetite of the investor. The case studies in Figures 7 and 8 demonstrate entropic hedging and pricing for call and digital options over a single investment horizon, and show the bid-offer spread around the mid model that prices the residual risk after hedging. To further develop intuition on the meaning of the hedge and price conditions and their solutions, it is useful to consider boundary cases where they simplify.
III.1 Empty market
In this section, assume that there are no underlying securities available to the investor for funding or hedging.
Assumption III.1 (Empty market).
The market of underlying securities is empty.
It is interesting to note that entropic pricing handles this trivial case in a sensible way. The calibration and hedge conditions are empty, and the price condition implies that the derivative price is a log-martingale in the expectation measure:
$p=\mathbb{E}[\alpha][P]$ (49)
In the absence of any market information, price reduces to expectation, adjusted to match the risk appetite of the investor.
III.2 Predictable funding
In this section, assume that the underlying market available to the investor includes a funding security whose final price is predictable.
Assumption III.2 (Predictable funding).
The market partitions into a vector of underlying securities with initial price $q$ and final price $q+dq$ and a funding security with strictly positive initial price $b>0$ and predictable final price $b+db$ that has zero variance $\mathbb{V}[db]=0$ in the expectation measure $\mathbb{E}$.
With this assumption, the calibration and hedge conditions are solved for the implied funding rate $r$ and implied funding spread $s$:
$r$ $=\frac{\mathbb{E}[db]}{b\,dt}$ (50)
$s$ $=0$
When no additional underlying securities are available, the price condition implies that the derivative price is a discounted log-martingale in the expectation measure:
$p=\frac{1}{1+r\,dt}\mathbb{E}[\alpha][P]$ (51)
More generally, underlying securities are available for hedging, and the calibration and hedge conditions for the unit optimal portfolio $\phi$ and hedge portfolio $\delta$ become:
00 0 $=\mathbb{E}^{\phi}[d\bar{q}]$ (52)
00 0 $=\mathbb{E}^{\phi}[d\bar{q}\,\exp[-\alpha(P-\delta\cdot d\bar{q})]]$
The price condition then derives the discounted log-martingale expression in the price measure:
$p=\frac{1}{1+r\,dt}\mathbb{E}^{\phi}[\alpha][P-\delta\cdot d\bar{q}]$ (53)
The funding security funds the investment and manifests as discounting in the price expression, and the additional underlying securities hedge the derivative risks.
III.3 Continuous diffusion
In this section, assume that the underlying and derivative prices diffuse continuously, and the investor is able to re-balance the hedge over short investment horizons.
Assumption III.3 (Continuous diffusion).
The diffusion of the underlying and derivative prices is continuous almost certainly in the expectation measure $\mathbb{E}$, so that their cumulant generating function satisfies:
$\log{}$ $\mathbb{E}\exp[\beta\,dp+\psi\cdot dq]=$ (54)
$\beta\mathbb{E}[dp]+\psi\cdot\mathbb{E}[dq]$
$+\frac{1}{2}\beta^{2}\mathbb{V}[dp]+\beta\psi\cdot\mathbb{V}[dq,dp]+\frac{1}{2}\psi\cdot\mathbb{V}[dq]\psi+O[dt^{2}]$
with $\mathbb{V}[dp]=O[dt]$, $\mathbb{V}[dq,dp]=O[dt]$ and $\mathbb{V}[dq]=O[dt]$.
Two assumptions are included here. The first assumption, that the market prices diffuse continuously, justifies the short-horizon expression for their cumulant generating function. The second assumption is that the investor re-balances frequently and at negligible cost, so that the $O[dt]$ corrections to the investment strategy can be ignored, and the bid-offer spread around the mid price derives exclusively from the post-hedging residual risks.
Moving to the price measure $\mathbb{E}^{\phi}$, the impact of the measure change on the cumulant generating function derives the drift adjustments in Girsanov’s theorem:
$\log{}$ $\mathbb{E}^{\phi}\exp[\beta\,dp+\psi\cdot dq]=$ (55)
$\beta(\mathbb{E}[dp]-\phi\cdot\mathbb{V}[dq,dp])+\psi\cdot(\mathbb{E}[dq]-\mathbb{V}[dq]\phi)$
$+\frac{1}{2}\beta^{2}\mathbb{V}[dp]+\beta\psi\cdot\mathbb{V}[dq,dp]+\frac{1}{2}\psi\cdot\mathbb{V}[dq]\psi+O[dt^{2}]$
Differentiating with respect to $\psi$ then generates:
$\frac{\displaystyle\mathbb{E}^{\phi}[dq\,\exp[\beta\,dp+\psi\cdot dq]]}{\displaystyle\mathbb{E}^{\phi}[\exp[\beta\,dp+\psi\cdot dq]]}=$ (56)
$\qquad\mathbb{E}[dq]+\beta\mathbb{V}[dq,dp]+\mathbb{V}[dq](\psi-\phi)+O[dt^{2}]$
These results are sufficient to solve the conditions. The calibration condition is solved for the implied funding rate $r$ and the unit optimal portfolio $\phi$:
$r$ $=\frac{\displaystyle\mathbb{V}[dq]^{-1}\mathbb{E}[dq]\cdot q}{\displaystyle\mathbb{V}[dq]^{-1}q\,dt\cdot q}+O[dt]$ (57)
$\phi$ $=\mathbb{V}[dq]^{-1}(\mathbb{E}[dq]-qr\,dt)+O[dt]$
and the hedge and price conditions are solved for the implied funding spread $s$, the hedge portfolio $\delta$ and the initial price $p$:
$s$ $=\frac{\displaystyle\mathbb{E}^{\phi}[P]-\mathbb{V}[dq]^{-1}\mathbb{V}[dq,P]\cdot q}{\displaystyle\mathbb{V}[dq]^{-1}q\,dt\cdot q}+O[dt]$ (58)
$\delta$ $=\mathbb{V}[dq]^{-1}(\mathbb{V}[dq,P]+qs\,dt)+O[dt]$
$p$ $=(1-r\,dt)\mathbb{E}^{\phi}[P]-\frac{1}{2}\alpha\mathbb{V}[P-\delta\cdot dq]+O[dt^{2}]$
III.4 Markovian diffusion
In this section, assume that the underlying and derivative prices are deterministic functions of a state vector.
Assumption III.4 (Markovian diffusion).
There is a vector $x$ of state variables whose cumulant generating function in the expectation measure $\mathbb{E}$ satisfies:
$\frac{1}{dt}\log{}$ $\mathbb{E}\exp[\eta\cdot dx]=$ (59)
$\eta\cdot\mu+\frac{1}{2}\eta\cdot\nu\eta+\int_{j}(\exp[\eta\cdot j]-1)\,\theta[dj]+O[dt]$
where the vector $\mu\equiv\mu[t,x]$ and positive-definite matrix $\nu\equiv\nu[t,x]$ are the mean and covariance rate for the continuous component of the diffusion, and the positive measure $\theta[dj]\equiv\theta[t,x][dj]$ is the frequency of jumps in the discontinuous component of the diffusion. The underlying price $q\equiv q[t,x]$ and derivative price $p\equiv p[t,x]$ are assumed to be functions of the time $t$ and state $x$.
This form for the cumulant generating function of the state implies its distribution is divisible, which enables the transition to the limit $dt\to 0$ of continuous hedging. Repeated differentiation with respect to $\eta$ generates Itô’s lemma for the function $f\equiv f[t,x]$ of the time $t$ and state $x$ in the expectation measure $\mathbb{E}$:
$\frac{1}{dt}\mathbb{E}[df]={}$ $\frac{\partial f}{\partial t}+\frac{\partial f}{\partial x}\cdot\mu+\frac{1}{2}\frac{\partial^{2}f}{\partial x^{2}}\cdot{}\!\cdot\nu$ (60)
$+\int_{j}\partial_{j}f\,\theta[dj]+O[dt]$
where $\partial_{j}f[t,x]:=f[t,x+j]-f[t,x]$. In the price measure $\mathbb{E}^{\phi}$, Itô’s lemma is modified by the drift adjustments of Girsanov’s theorem:
$\frac{1}{dt}\mathbb{E}^{\phi}[df]={}$ $\frac{\partial f}{\partial t}+\frac{\partial f}{\partial x}\cdot\mu^{\phi}+\frac{1}{2}\frac{\partial^{2}f}{\partial x^{2}}\cdot{}\!\cdot\nu$ (61)
$+\int_{j}\partial_{j}f\,\theta^{\phi}[dj]+O[dt]$
where the mean rate and frequency of jumps are adjusted:
$\mu^{\phi}$ $:=\mu-\nu\left(\phi\cdot\frac{\partial q}{\partial x}\right)$ (62)
$\theta^{\phi}[dj]$ $:=\exp[-\phi\cdot\partial_{j}q]\,\theta[dj]$
With this assumption, the price model is established on the assumed divisible process for the state vector $x$ and the underlying price vector $q[t,x]$ given as a function of time and state. Using the expression for Itô’s lemma in the price measure, the calibration, hedge and price conditions are transformed into partial differential-integral form. The calibration condition becomes:
$0={}$ $\frac{\partial q}{\partial t}-qr+\frac{\partial q}{\partial x}\cdot\mu^{\phi}+\frac{1}{2}\frac{\partial^{2}q}{\partial x^{2}}\cdot{}\!\cdot\nu$ (63)
$+\int_{j}\partial_{j}q\,\theta^{\phi}[dj]+O[dt]$
which is solved simultaneously with the self-funding condition $\phi\cdot q=0$ for the implied funding rate $r$ and the unit optimal portfolio $\phi$. The hedge and price conditions become:
$0={}$ $\frac{\partial p}{\partial x}\cdot\nu\frac{\partial q}{\partial x}+qs-\left(\frac{\partial q}{\partial x}\cdot\nu\frac{\partial q}{\partial x}\right)\delta$ (64)
$\!\!\!\!+\int_{j}\frac{1}{\alpha}(1-\exp[-\alpha(\partial_{j}p-\delta\cdot\partial_{j}q)])\,\partial_{j}q\,\theta^{\phi}[dj]+O[dt]$
$0={}$ $\frac{\partial p}{\partial t}-pr+\frac{\partial p}{\partial x}\cdot\mu^{\phi}+\frac{1}{2}\frac{\partial^{2}p}{\partial x^{2}}\cdot{}\!\cdot\nu$
$-\frac{1}{2}\alpha\left(\frac{\partial p}{\partial x}-\delta\cdot\frac{\partial q}{\partial x}\right)\cdot\nu\left(\frac{\partial p}{\partial x}-\delta\cdot\frac{\partial q}{\partial x}\right)$
$+\int_{j}\frac{1}{\alpha}(1-\exp[-\alpha(\partial_{j}p-\delta\cdot\partial_{j}q)])\,\theta^{\phi}[dj]+O[dt]$
which are solved simultaneously with the self-funding condition $\delta\cdot q=p$ for the implied funding spread $s$, hedge portfolio $\delta$ and derivative price $p$.
Starting with its terminal contractual settlement, the numerical solution of these conditions for the derivative price repeats a Newton-Raphson scheme that determines the initial price $p$ from the final price $p+dp$ over each interval in a discretised schedule. The high dimensional complexity of this procedure is a consequence of the non-linearity of the hedge portfolio in the hedge condition. There are two useful scenarios where this condition is linear: the mid-pricing case $\alpha=0$; and the continuous diffusion case $\theta[dj]=0$. Pricing simplifies in these cases.
Mid-pricing
When $\alpha=0$, the hedge and price conditions simplify:
$s={}$ $\frac{\displaystyle p-V^{-1}\left(\frac{\partial p}{\partial x}\cdot\nu\frac{\partial q}{\partial x}+\int_{j}\partial_{j}p\,\partial_{j}q\,\theta^{\phi}[dj]\right)\cdot q}{\displaystyle V^{-1}q\cdot q}+O[dt]$
$\delta={}$ $V^{-1}\left(\frac{\partial p}{\partial x}\cdot\nu\frac{\partial q}{\partial x}+\int_{j}\partial_{j}p\,\partial_{j}q\,\theta^{\phi}[dj]+qs\right)+O[dt]$
$0={}$ $\frac{\partial p}{\partial t}-pr+\frac{\partial p}{\partial x}\cdot\mu^{\phi}+\frac{1}{2}\frac{\partial^{2}p}{\partial x^{2}}\cdot{}\!\cdot\nu$ (65)
$+\int_{j}(\partial_{j}p-\delta\cdot\partial_{j}q)\,\theta^{\phi}[dj]+O[dt]$
where:
$V$ $:=\frac{\partial q}{\partial x}\cdot\nu\frac{\partial q}{\partial x}+\int_{j}\partial_{j}q\,\partial_{j}q\,\theta^{\phi}[dj]$ (66)
Embedding the expressions for the implied funding spread and hedge portfolio, the price condition becomes a linear parabolic differential-integral equation for the derivative price. This equation balances the contributions to the return from the continuous and discontinuous components of the diffusion. The underlying market is incomplete and the hedge strategy does not eliminate market risk, but this choice for the hedge portfolio optimises the return after hedging.
Continuous diffusion
When $\theta[dj]=0$, the hedge and price conditions simplify:
$s={}$ $\frac{\displaystyle p-V^{-1}\left(\frac{\partial p}{\partial x}\cdot\nu\frac{\partial q}{\partial x}\right)\cdot q}{\displaystyle V^{-1}q\cdot q}+O[dt]$
$\delta={}$ $V^{-1}\left(\frac{\partial p}{\partial x}\cdot\nu\frac{\partial q}{\partial x}+qs\right)+O[dt]$
$0={}$ $\frac{\partial p}{\partial t}-pr+\frac{\partial p}{\partial x}\cdot\mu^{\phi}+\frac{1}{2}\frac{\partial^{2}p}{\partial x^{2}}\cdot{}\!\cdot\nu$ (67)
$-\frac{1}{2}\alpha\left(\frac{\partial p}{\partial x}-\delta\cdot\frac{\partial q}{\partial x}\right)\cdot\nu\left(\frac{\partial p}{\partial x}-\delta\cdot\frac{\partial q}{\partial x}\right)+O[dt]$
where:
$V$ $:=\frac{\partial q}{\partial x}\cdot\nu\frac{\partial q}{\partial x}$ (68)
Embedding the expressions for the implied funding spread and hedge portfolio, the price condition becomes a non-linear parabolic differential equation for the derivative price. This equation modifies the expression for the price to adjust for residual risk. The underlying market is incomplete and the hedge strategy does not eliminate market risk, but this choice for the hedge portfolio optimises the return after hedging.
III.5 Heston model
In this section, the Heston model is proposed for the expectation measure. This concrete example of a continuous Markovian diffusion illustrates the key features of the more general framework.
Assumption III.5 (Heston model).
The state comprises a funding price $b$, an underlying price $q$ and a variance $\nu$ whose cumulant generating function in the expectation measure $\mathbb{E}$ satisfies:
$\frac{1}{dt}\log{}$ $\mathbb{E}\exp[\beta\,db+\psi\,dq+\eta\,d\nu]=$ (69)
$\beta\,br+\psi\,q\mu+\eta\,(\bar{\nu}-\nu)\kappa$
$+\frac{1}{2}\psi^{2}\,q^{2}\nu+\psi\eta\,q\nu\rho\sigma+\frac{1}{2}\eta^{2}\,\nu\sigma^{2}+O[dt]$
The funding price has deterministic increment with mean $br\,dt$ where $r$ is the funding rate. The underlying price has increment with mean $q\mu\,dt$ and variance $q^{2}\nu\,dt$ where $\mu$ is the mean rate. The variance has increment with mean $(\bar{\nu}-\nu)\kappa\,dt$ and variance $\nu\sigma^{2}\,dt$ where $\kappa$ is the mean reversion rate, $\bar{\nu}$ is the mean reversion level and $\sigma$ is the volatility of variance. The final parameter $\rho$ is the correlation between the increments of the underlying price and the variance.
Since the funding security has deterministic return, the implied funding rate is $r$ and the implied funding spread is $s=0$. In order to satisfy the self-funding condition, the portfolio with underlying weight $\phi$ must have funding weight $-\phi q/b$, and the cumulant generating function in the price measure $\mathbb{E}^{\phi}$ satisfies:
$\frac{1}{dt}\log{}$ $\mathbb{E}^{\phi}\exp[\beta\,db+\psi\,dq+\eta\,d\nu]=$ (70)
$\beta\,br+\psi\,q\mu^{\phi}+\eta\,(\bar{\nu}^{\phi}-\nu)\kappa$
$+\frac{1}{2}\psi^{2}\,q^{2}\nu+\psi\eta\,q\nu\rho\sigma+\frac{1}{2}\eta^{2}\,\nu\sigma^{2}+O[dt]$
where the mean rate of the underlying price and the mean reversion level of the variance are adjusted:
$\mu^{\phi}$ $:=\mu-\phi q\nu$ (71)
$\bar{\nu}^{\phi}$ $:=\bar{\nu}-\phi q\nu\frac{\rho\sigma}{\kappa}$
The calibration condition then implies that the unit optimal portfolio has $\phi=(\mu-r)/(q\nu)$, so that:
$\mu^{\phi}$ $=r$ (72)
$\bar{\nu}^{\phi}$ $=\bar{\nu}-(\mu-r)\frac{\rho\sigma}{\kappa}$
The state is also a Heston model in the price measure, albeit with these two drift parameters adjusted to account for the change of measure. In this approach, the Heston parameters in the price measure can be estimated from statistical analysis of the state in the expectation measure, applying the measure adjustment post-estimation to calibrate to the underlying price.
The hedge and price conditions in the Heston model generate the expressions:
$\delta={}$ $\frac{\partial p}{\partial q}+\frac{\partial p}{\partial\nu}\frac{\rho\sigma}{q}+O[dt]$ (73)
$0={}$ $\frac{\partial p}{\partial t}-pr+\frac{\partial p}{\partial b}br+\frac{\partial p}{\partial q}qr+\frac{\partial p}{\partial\nu}((\bar{\nu}-\nu)\kappa-(\mu-r)\rho\sigma)$
$+\frac{1}{2}\frac{\partial^{2}p}{\partial q^{2}}\,q^{2}\nu+\frac{\partial^{2}p}{\partial q\,\partial\nu}\,q\nu\rho\sigma+\frac{1}{2}\frac{\partial^{2}p}{\partial\nu^{2}}\,\nu\sigma^{2}$
$-\frac{1}{2}\alpha\left(\frac{\partial p}{\partial\nu}\right)^{\!\!2}\!\nu\sigma^{2}(1-\rho^{2})+O[dt]$
for the hedge portfolio $\delta$ and the derivative price $p$. In the first expression, the hedge ratio is adjusted when the increment of the variance is correlated with the increment of the underlying price. The derivative price then satisfies a linear parabolic differential equation with a non-linear correction term that prices the unhedgeable risks when the increments are not perfectly correlated.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 9: In these examples of data-driven models, the underlying payoff $Q:=q+dq$ is modelled in the expectation measure as a uniform variable on 100 samples from a standard normal variable. In each row, the price measure is supported on the same samples, adjusting the probabilities to match the target underlying price $q$ with minimum entropy relative to the expectation measure. No adjustment is necessary when $q=0$, and the probabilities are downward/upward sloping when $q$ is negative/positive, biasing the distribution in the direction of the target underlying price.
      The graphs plot the delta and price per unit notional as a function of the strike $k$ of a call option with unit payoff $P:=(Q-k)^{+}$ for a range of notionals $\lambda$, highlighting the central plot for the mid-pricing case $\lambda=0$. Risk aversion is fixed at one, $\alpha=1$, and the implied funding rate is set to zero, $r=0$.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 10: In these examples of data-driven models, the underlying payoff $Q:=q+dq$ is modelled in the expectation measure as a uniform variable on 100 samples from a standard normal variable. In each row, the price measure is supported on the same samples, adjusting the probabilities to match the target underlying price $q$ with minimum entropy relative to the expectation measure. No adjustment is necessary when $q=0$, and the probabilities are downward/upward sloping when $q$ is negative/positive, biasing the distribution in the direction of the target underlying price.
      The graphs plot the delta and price per unit notional as a function of the strike $k$ of a digital option with unit payoff $P:=1_{Q>k}$ for a range of notionals $\lambda$, highlighting the central plot for the mid-pricing case $\lambda=0$. Risk aversion is fixed at one, $\alpha=1$, and the implied funding rate is set to zero, $r=0$.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 11: In these examples of data-driven models, the underlying payoff $Q:=q+dq$ is modelled in the expectation measure as a uniform variable on 10,000 samples from a standard normal variable. In each row, the price measure is supported on the same samples, adjusting the probabilities to match the target underlying price $q$ with minimum entropy relative to the expectation measure. No adjustment is necessary when $q=0$, and the probabilities are downward/upward sloping when $q$ is negative/positive, biasing the distribution in the direction of the target underlying price.
      The graphs plot the delta and price per unit notional as a function of the strike $k$ of a call option with unit payoff $P:=(Q-k)^{+}$ for a range of notionals $\lambda$, highlighting the central plot for the mid-pricing case $\lambda=0$. Risk aversion is fixed at one, $\alpha=1$, and the implied funding rate is set to zero, $r=0$.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 12: In these examples of data-driven models, the underlying payoff $Q:=q+dq$ is modelled in the expectation measure as a uniform variable on 10,000 samples from a standard normal variable. In each row, the price measure is supported on the same samples, adjusting the probabilities to match the target underlying price $q$ with minimum entropy relative to the expectation measure. No adjustment is necessary when $q=0$, and the probabilities are downward/upward sloping when $q$ is negative/positive, biasing the distribution in the direction of the target underlying price.
      The graphs plot the delta and price per unit notional as a function of the strike $k$ of a digital option with unit payoff $P:=1_{Q>k}$ for a range of notionals $\lambda$, highlighting the central plot for the mid-pricing case $\lambda=0$. Risk aversion is fixed at one, $\alpha=1$, and the implied funding rate is set to zero, $r=0$.
III.6 Data-driven model
In this section, assume that the model is supported on a finite set of states with uniform probability in the expectation measure.
Assumption III.6 (Data-driven model).
The model is supported on a finite set of states that are equally likely in the expectation measure. In the $n$th state, the underlying return is $dq_{n}$ and the final derivative price is $P_{n}$.
For this data-driven approach, the expectation measure is represented as a uniform distribution on a finite set of states, which could for example be sampled from a generative model trained on historical market data. The calibration condition is solved for the unit optimal portfolio $\phi$ via the optimisation problem:
$\phi=\operatorname*{arg\,min}\nolimits_{\psi}\!\left\{\sum\nolimits_{n}\exp[-\psi\cdot dq_{n}]:\psi\cdot q=0\right\}$ (74)
The price measure is then supported on the same set of states but with non-uniform probabilities:
$\gamma_{n}:=\frac{\displaystyle\exp[-\phi\cdot dq_{n}]}{\displaystyle\sum\nolimits_{m}\exp[-\phi\cdot dq_{m}]}$ (75)
The hedge and price conditions are solved simultaneously for the hedge portfolio $\delta$ and the initial derivative price $p$ via the optimisation problem:
$\delta$ $=\operatorname*{arg\,min}\nolimits_{\epsilon}\!\left\{\sum\nolimits_{n}\gamma_{n}\exp[-\alpha(P_{n}-\epsilon\cdot dq_{n})]:\epsilon\cdot q=p\right\}$
$p$ $=-\frac{1}{\alpha}\log\sum\nolimits_{n}\gamma_{n}\exp[-\alpha(P_{n}-\delta\cdot dq_{n})]$ (76)
These expressions are amenable to machine learning methods and are the starting point for a variety of applications in deep hedging.
This example demonstrates the reach of the entropic framework. With no theory to support development beyond the information that the data expresses, data-driven models need a more general foundation that does not depend on unrealistic assumptions such as complete markets, continuous hedging or risk-free funding. Advocates of deep hedging address this by replacing dynamic risk neutralisation with dynamic risk optimisation, using convex risk metrics to assess strategic performance. While they suffer from lack of explainability, data-driven models can uncover complex relationships between economic variables that may be hard to capture theoretically.
IV Market and model risk
Returning to the general pricing framework, with the expectation measure $\mathbb{E}$ the calibration conditions:
00 0 $=\phi\cdot q$ (77)
00 0 $=\mathbb{E}^{\phi}[dq]-qr\,dt$
are solved to generate the implied funding rate $r$ and the unit optimal portfolio $\phi$ for the underlying securities, and the hedge and price conditions:
$p$ $=\delta\cdot q$ (78)
00 0 $=\mathbb{E}^{\phi}[(dq-q(r+\alpha s)\,dt)\exp[-\alpha(P-\delta\cdot dq)]]$
$p$ $=\mathbb{E}^{\phi}[\alpha][P-\delta\cdot dq]$
are solved to generate the implied funding spread $s$, the hedge portfolio $\delta$ and the initial price $p$ for the derivative security. While the hedge neutralises the underlying component of the derivative risk, when the market is incomplete there are unhedged risks that are sensitive to the choice of expectation measure. The entropic pricing framework thus distinguishes between the market and model risks in the derivative security.
Consider an alternative choice $\bar{\mathbb{E}}$ for the expectation measure, related to the original expectation measure $\mathbb{E}$ by the Radon-Nikodym weight:
$\frac{d\bar{\mathbb{E}}}{d\mathbb{E}}=\frac{\displaystyle\exp[\epsilon W]}{\displaystyle\mathbb{E}[\exp[\epsilon W]]}=1+\epsilon(W-\mathbb{E}[W])+O[\epsilon^{2}]$ (79)
for a bounded weight $W$ and a scale parameter $\epsilon$. Using this alternative expectation measure in the calibration conditions, the implied funding rate and unit optimal portfolio are adjusted, $\bar{r}=r+\epsilon\,\partial r$ and $\bar{\phi}=\phi+\epsilon\,\partial\phi$, with:
$\partial r$ $=\frac{\displaystyle\mathbb{V}^{\phi}[dq]^{-1}\mathbb{V}^{\phi}[dq,W]\cdot q}{\displaystyle\mathbb{V}^{\phi}[dq]^{-1}q\,dt\cdot q}+O[\epsilon]$ (80)
$\partial\phi$ $=\mathbb{V}^{\phi}[dq]^{-1}(\mathbb{V}^{\phi}[dq,W]-q\bar{r}\,dt)+O[\epsilon]$
Applying these adjustments, the original and alternative price measures are then related by:
$\bar{\mathbb{E}}^{\bar{\phi}}[X]-\mathbb{E}^{\phi}[X]=$ (81)
$\qquad\epsilon\mathbb{V}^{\phi}[W-\partial\phi\cdot dq,X]+O[\epsilon^{2}]$
$\bar{\mathbb{E}}^{\bar{\phi}}[\alpha][X]-\mathbb{E}^{\phi}[\alpha][X]=$
$\qquad\epsilon\mathbb{V}^{\phi}\!\left[W-\partial\phi\cdot dq,-\frac{1}{\alpha}\frac{\displaystyle\exp[-\alpha X]}{\displaystyle\mathbb{E}^{\phi}[\exp[-\alpha X]]}\right]+O[\epsilon^{2}]$
for the undetermined variable $X$.
Similar adjustments apply to the implied funding spread, hedge portfolio and initial price, derived from the hedge and price conditions in the alternative expectation measure:
$\bar{p}$ $=\bar{\delta}\cdot q$ (82)
00 0 $=\bar{\mathbb{E}}^{\bar{\phi}}[(dq-q(\bar{r}+\alpha\bar{s})\,dt)\exp[-\alpha(P-\bar{\delta}\cdot dq)]]$
$\bar{p}$ $=\bar{\mathbb{E}}^{\bar{\phi}}[\alpha][P-\bar{\delta}\cdot dq]$
Focusing on the price condition, the adjustment to the initial price is $\bar{p}=p+\epsilon\,\partial p$ with:
$\partial p=\frac{1}{1+(r+\alpha s)\,dt}\times$ (83)
$\quad\mathbb{V}^{\phi}\!\left[W-\partial\phi\cdot dq,-\frac{1}{\alpha}\frac{\displaystyle\exp[-\alpha(P-\delta\cdot dq)]}{\displaystyle\mathbb{E}^{\phi}[\exp[-\alpha(P-\delta\cdot dq)]]}\right]+O[\epsilon]$
The hedge portfolio is the metric of market risk in the entropic framework, and this adjustment is the metric of model risk as it measures the sensitivity of the initial price to the choice of expectation measure. If the underlying market is complete, the variance of the hedged return is zero and the sensitivity to the expectation measure is eliminated. More generally, model risk increases in proportion with the degree of market incompleteness.
V Entropic XVA
An unconstrained strategy that maximises the mean does not incentivise risk management and results in a portfolio with potentially disastrous downside risks. For this reason, a strategy that instead aims to maximise the risk-adjusted mean creates a more balanced portfolio that can be fine-tuned to the risk appetite of the investor.
It would appear from this statement that risk management is an investment choice, with the risk appetite a free parameter under the control of the investor. This is not the case. Without risk management there is a material risk that the investor will not meet their contractual commitments. At the very least, this potentially tarnishes the reputation of the investor as a reliable counterparty, with consequent impact on the prices they will be offered. In regulated markets, risk management is also a legal requirement with detailed and specific obligations for margin and capital.
Figure 13: In this example, the trader receives the initial derivative price $p$ and deploys additional capital $c$ to guarantee settlement, with $c+p\geq 0$. These proceeds are deposited in a margin account with unit price $b>0$. The final derivative price $p+dp$ is settled up to the value $(c+p)(1+db/b)$ of the margin account, with any remainder returned to the trader. Default occurs when the margin account is insufficient to meet the contractual obligation. In the expression for the final settlement, the cap is denoted by the superscript.
Figure 14: Two counterparties deploy capital $c_{1}$ and $c_{2}$ respectively as their guarantee of settlement, with $c_{1}-p\geq 0$ and $c_{2}+p\geq 0$, and they default on their contractual obligation if this is insufficient. Net of all considerations, the counterparties share the combined proceeds from the margin accounts, and the final return on capital is floored and capped by the respective default events. In the expressions for the final settlement and capital, the floor is denoted by the subscript and the cap is denoted by the superscript.
Figure 15: Eliminating internal settlements, the trade is simplified as a funding pool with contributions from both counterparties shared according to the contractual settlement of the derivative security adjusted for default. Using the entropic margin model, these settlements can be encoded in a smart contract for decentralised distribution.
If the true costs of risk management are fully accounted in the return, reckless investments are naturally penalised by their associated margin costs. Revisiting the bilateral model outlined in the introduction, suppose that two counterparties deploy capital $c_{1}$ and $c_{2}$ to a trade with initial price $p$. Excess capital is deposited in the margin account with unit price $b>0$. The trade is settled from the margin account with no additional funds injected, creating the risk that insufficient capital results in default by one of the counterparties.
As shown in Figures 15, 15 and 15, net of all considerations the capital returned from the trade is:
$c_{1}+dc_{1}$ $=(b+db)\!\left(\frac{c_{1}}{b}+d\frac{p}{b}\right)_{0}^{(c_{1}+c_{2})/b}$ (84)
$=c_{1}(1+db/b)+\left(dp-\frac{p}{b}db\right)_{-c_{1}(1+db/b)}^{c_{2}(1+db/b)}$
$c_{2}+dc_{2}$ $=(b+db)\!\left(\frac{c_{2}}{b}-d\frac{p}{b}\right)_{0}^{(c_{1}+c_{2})/b}$
$=c_{2}(1+db/b)-\left(dp-\frac{p}{b}db\right)_{-c_{1}(1+db/b)}^{c_{2}(1+db/b)}$
where the subscript denotes the floor and the superscript denotes the cap. The combined proceeds from the margin accounts are shared between the counterparties, adjusting for the funded return on the derivative security, floored and capped so that the capital is positive.
The first counterparty defaults when the floor is breached and the second counterparty defaults when the cap is breached. Each counterparty assesses the threshold for default using their expectation measure $\mathbb{E}_{1}$ and $\mathbb{E}_{2}$, and default is avoided when the capital exceeds the maximum liability:
$c_{1}$ $\geq p-\operatorname*{ess\,inf}\nolimits_{1}\!\left[\frac{P}{1+db/b}\right]$ (85)
$c_{2}$ $\geq\operatorname*{ess\,sup}\nolimits_{2}\!\left[\frac{P}{1+db/b}\right]-p$
where $P:=p+dp$ is the final price. Setting these inequalities to equalities generates the minimum capital required to fully guarantee settlement, under the assumptions of the respective expectation measures. More generally, the counterparties can exchange certainty of settlement for a reduction in capital. In the entropic margin model, this is parametrised by the default aversion $\beta_{1}$ and $\beta_{2}$ with:
$m_{1}$ $:=c_{1}-p=-\mathbb{E}_{1}[\beta_{1}]\!\left[\frac{P}{1+db/b}\right]$ (86)
$m_{2}$ $:=c_{2}+p=\mathbb{E}_{2}[-\beta_{2}]\!\left[\frac{P}{1+db/b}\right]$
where the case $\beta_{1}=\infty$ means the first counterparty provides full default protection and the case $\beta_{2}=\infty$ means the second counterparty provides full default protection.
Viability of the trade is assessed by the counterparty using their risk aversion $\alpha_{1}$ and $\alpha_{2}$. Consider the threshold for execution given by the condition:
$\mathbb{E}_{1}[\alpha_{1}]\!\left[\frac{c_{1}+dc_{1}}{1+db/b}\right]$ $\geq c_{1}$ (87)
$\mathbb{E}_{2}[\alpha_{2}]\!\left[\frac{c_{2}+dc_{2}}{1+db/b}\right]$ $\geq c_{2}$
on the entropy-adjusted mean return on funded capital. These conditions impose bounds on the viable price:
$p$ $\leq p_{1}:=\mathbb{E}_{1}[\alpha_{1}]\!\left[\left(\frac{P}{1+db/b}\right)_{\!-m_{1}}^{\!m_{2}}\right]$ (88)
$p$ $\geq p_{2}:=\mathbb{E}_{2}[-\alpha_{2}]\!\left[\left(\frac{P}{1+db/b}\right)_{\!-m_{1}}^{\!m_{2}}\right]$
The trade is viable when the initial price $p$ is in the range $p_{2}\leq p\leq p_{1}$, and is therefore not viable when $p_{1}<p_{2}$. Price is thus constrained to the sector bounded by these viability thresholds, equilibrating to a price within the range. Since the entropy-adjusted mean is monotonic as a function of the notional, the range narrows as the notional increases and there is an equilibrium point at which no further trades are viable. Equilibrium is thus discovered as the solution of the relationship:
[…]
for the initial price per unit notional $p$ and the equilibrium notional $\lambda$. Terminating when the trade exceeds the risk appetite of the investor, this procedure balances the differing expectations of the counterparties to identify their equilibrium price, accounting for their respective funding costs and risks of default.
Refer to caption Refer to caption
$\lambda=0$ and $c=\infty$ $\lambda=0$ and $c=0.5$
Refer to caption Refer to caption
$\lambda=1$ and $c=\infty$ $\lambda=1$ and $c=0.5$
Refer to caption Refer to caption
$\lambda=2$ and $c=\infty$ $\lambda=2$ and $c=0.5$
Figure 16: In this example, two counterparties have differing expectations for the future state. The first counterparty models the state as a uniform variable with zero mean and unit standard deviation. The second counterparty models the state as a normal variable with mean $\mu$ and unit standard deviation. The trade is then a call option on the state with zero strike. In each graph, the entropy-adjusted mean return on capital, computed with risk aversion $\alpha=1$, is plotted for both counterparties as a function of the option price per unit notional, including a range of values for the mean $\mu$ assumed by the second counterparty. The first counterparty is fully capitalised against default, and each row includes cases $c=\infty,0.5$ for the capital of the second counterparty. Each column includes cases $\lambda=0,1,2$ for the notional, where the case $\lambda=0$ corresponds to mid-pricing.
      The trade is viable when the entropy-adjusted mean return on capital is positive for both counterparties. The first counterparty buys the option when its price is low, and the second counterparty sells the option when its price is high. The second counterparty is more likely to sell when the mean $\mu$ is lower, as this reduces the expected option payoff. Model risk increases as the notional $\lambda$ is increased, which pushes the threshold prices in opposite directions and makes the trade less viable. Finally, for the second counterparty, decreasing the capital $c$ increases the risk of default. The CVA cost to the first counterparty and DVA benefit to the second counterparty then pushes both threshold prices lower.
VI Literature review
Understanding the relationship between the economic and price measures has been the fundamental question of mathematical finance ever since Bachelier first applied stochastic calculus in his pioneering thesis [1]. In this thesis, Bachelier develops a model for the logarithm of the security price as Brownian motion around its equilibrium, deriving expressions for options that would not look unfamiliar today. Methods of portfolio optimisation originated in the work of Markowitz [2] and Sharpe [3] using Gaussian statistics for the returns, generalised by later authors to allow more sophisticated distributional assumptions and measures of utility. These approaches are embedded in the economic measure, explaining the origin of price as the equilibrium of market activity uncovering the expectations of participants.
The impact of dynamic hedging on price was first recognised in the articles by Black and Scholes [4] and Merton [5]. Adopting the framework devised by Bachelier, these authors observe that the option return is replicated exactly by a strategy that continuously offsets the delta of the option price to the underlying price. The theory matured with the work of Harrison and Pliska [6, 7], providing a precise statement of the conditions for market completeness and the martingale property of price in a measure equivalent to the economic measure. The discipline has since expanded in numerous directions, with significant advances for term structure and default modelling and numerical methods for complex derivative structures.
Closer consideration of market incompleteness led to the development of alternative pricing principles based on the dynamic optimisation of convex risk metrics, as presented in [8, 9, 10]. With foundations that extend well beyond the domain of applicability of risk-neutral pricing, these principles enabled the development of data-driven methods including the pioneering work on deep hedging by Buehler and others [11, 12]. As a recent practical example, Oya [13] applies these methods to explain the mark-to-market valuation adjustment for Bermudan swaptions.
Entropy methods have found application across the range of mathematical finance, including portfolio optimisation [14] and derivative pricing [15, 16] – see also the review essay [17] which includes further references. The article [18] by Frittelli proposes the minimal relative entropy measure as a solution to the problem of pricing in incomplete markets, and links this solution with the maximisation of expected exponential utility. The entropic solution is connected with risk optimisation in the introduction [19], which this essay extends with the adoption of the entropic risk metric. As a measure of disorder, entropy performs a role similar to variance but is better suited to the real distributions of market returns. Proponents of the use of entropy justify the approach by appeal to the modelling of information flows in dynamical systems and the analogy with thermodynamics.
While market evidence for discounting basis existed earlier, the basis widening that occurred as a result of the Global Financial Crisis of 2007-2008 motivated research into funding and its impact on discounting. Work by Johannes and Sundaresan [20], Fujii and Takahashi [21], Piterbarg and Antonov [22, 23, 24], Henrard [25], McCloud [26] and others established its theoretical justification and practical application.
Most approaches to derivative pricing begin with the assumption of continuous settlement, and stochastic calculus is an essential ingredient for these developments. The representation of the stochastic differential equation used in this article follows the discoveries of Lévy [27], Khintchine [28] and Itô [29]. The technical requirements of the Lévy-Khintchine representation can be found in these articles and other standard texts in probability theory. The change from economic measure to price measure implied by the entropic principle extends the result from Girsanov [30] to include the scaling adjustment of the jump frequency in addition to the drift adjustment. The example model referenced here is the Heston model [31], a popular model in quantitative finance.
It is remarkable that the principle of dynamic convex risk optimisation enables the consistent and simultaneous application of models across electronic trading algorithms, for pricing and hedging exotic derivatives, and to determine economic and regulatory capital. In this context, the entropic risk metric is promoted for its simple implementation as an extension of the risk-neutral framework, which supports integration into legacy platforms.